# Hype, Sustainability, and the Price of the Bigger-is-Better Paradigm in AI

### v1

@@ 6 Ways Forward: small can also be beautiful
Through defining our research agendas (which papers and talks are accepted, which questions are asked), the research community can shape how AI progresses and is appreciated. We can deepen our understanding of performance by diversifying our benchmarks, investigating the limitations of these benchmarks, and building bridges to other communities to pursue tasks and problems that are relevant in different contexts. All work on large AI system should also be compared to simpler baselines, to open the conversation on trade-offs between scale and other factors. And we should value research on open research questions such as uncertainty quantification or causality, even when it is conducted on smaller models, as it can bring valuable insights for the field as a whole.
@@ A.2 Imputation of missing values
Not all features of notable AI systems are reported, for instance inference cost, compute cost in dollars, may often be missing. For all plots but the one about dataset size [^3] We did not impute the dataset size as, unlike the other quantity, it does not need to have links to other quantities. we impute the missing quantities from the reported one. Specifically, we use an imputation with chained equations (Van Buuren & Groothuis-Oudshoorn, 2011, MICE). We relate number of parameters, training FLOP, inference FLOP, training cost, to each other as well as to dataset size, number of epochs, application domain, and time the model was published.
@@ Appendix C Labor and the human cost
While we, as a computer-science community, are well attuned to computational costs, other costs come in when developing large AI models. In particular, recent breakthroughs in ML rely on significant human labor to create data at various points throughout the development and deployment life cycle. ImageNet, the dataset that helped catalyze the current AI era, was only possible due to significant click worker labor, namely Amazon Mechanical Turk, which replaced costly and slow student labor (Hempel, 2018; Deng et al., 2009). Subsequently, the Mechanical Turk platform, and the ability it gave companies and researchers to remotely direct large numbers of low paid workers, became critical to construct emblematic computer vision datasets like MS-COCO (Lin et al., 2014) and Flickr-30K (Plummer et al., 2015). More recently, reinforcement learning from human feedback (RLHF) has become a central to conversational large language models, allowing them to better respond to human preferences (Ziegler et al., 2020) and to provide outputs that stay within the boundaries of politically and socially acceptable expression. In fact, RLHF was hailed as the ’secret ingredient’ of ChatGPT (Metz, 2023). RLHF continues the legacy of Mechanical Turk, paying workers, often in developing countries, a few dollars an hour for labor that is repetitive, exhausting, and often emotionally taxing (Hao & Seetharaman, 2023). Even though each worker is paid little, given the scale of RLHF endeavors, it remains inaccessible to average research labs and startups, who cannot afford to pay the thousands of people required to carry out this tuning (Widder et al., 2023).

### v2

@@ 6 Ways Forward: small can also be beautiful
By defining research agendas (which papers and talks are accepted, which questions are asked), the research community can shape how AI progresses and is appreciated. We can deepen our understanding of performance by diversifying our benchmarks, investigating the limitations of these benchmarks, and building bridges to other communities to pursue tasks and problems that are relevant in different contexts. All work on large AI system should also be compared to simpler baselines, to open the conversation on trade-offs between scale and other factors. And we should value research on open research questions such as uncertainty quantification or causality, even when it is conducted on smaller models, as it can bring valuable insights for the field as a whole.
@@ A.2 Imputation of missing values
Not all features of notable AI systems are reported, for instance inference cost, compute cost in dollars, may often be missing. For all plots but the one about dataset size [^3] We did not impute the dataset size as, unlike the other quantity, it does not need to have links to other quantities. we impute the missing quantities from the reported one. Specifically, we use an imputation with chained equations (Van Buuren and Groothuis-Oudshoorn, 2011, MICE). We relate number of parameters, training FLOP, inference FLOP, training cost, to each other as well as to dataset size, number of epochs, application domain, and time the model was published.
@@ Appendix C Labor and the human cost
While we, as a computer-science community, are well attuned to computational costs, other costs come in when developing large AI models. In particular, recent breakthroughs in ML rely on significant human labor to create data at various points throughout the development and deployment life cycle. ImageNet, the dataset that helped catalyze the current AI era, was only possible due to significant click worker labor, namely Amazon Mechanical Turk, which replaced costly and slow student labor (Hempel, 2018; Deng et al., 2009). Subsequently, the Mechanical Turk platform, and the ability it gave companies and researchers to remotely direct large numbers of low paid workers, became critical to construct emblematic computer vision datasets like MS-COCO (Lin et al., 2014) and Flickr-30K (Plummer et al., 2015). More recently, reinforcement learning from human feedback (RLHF) has become a central to conversational large language models, allowing them to better respond to human preferences (Ziegler et al., 2020) and to provide outputs that stay within the boundaries of politically and socially acceptable expression. In fact, RLHF was hailed as the ’secret ingredient’ of ChatGPT (Metz, 2023). RLHF continues the legacy of Mechanical Turk, paying workers, often in developing countries, a few dollars an hour for labor that is repetitive, exhausting, and often emotionally taxing (Hao and Seetharaman, 2023). Even though each worker is paid little, given the scale of RLHF endeavors, it remains inaccessible to average research labs and startups, who cannot afford to pay the thousands of people required to carry out this tuning (Widder et al., 2023).